Find the equation of the ellipse when foci is (5,0) and (-5,0) and the two vertices are (13,0) and (-13,0)
Answer by lwsshak3(11628)

Find the equation of the ellipse when foci is (5,0) and (-5,0) and the two vertices are (13,0) and (-13,0)
**
Standard form of ellipse with horizontal major axis: (x-h)^2/a^2+(y-k)^2/b^2=1, a>b, (h,k) being the (x,y) coordinates of the center.
For given equation for ellipse:
Center: (0,0)
length of horizontal major axis=26=2a
a=13
a^2=169
..
c=5
c^2=25.
..
c^2=a^2-b^2
b^2=a^2-c^2=169-25=144
b=12
..
Equation of given ellipse:
(x-0)^2/25+(y-0)^2=1
x^2/169+y^2/144=1